# FreqPolicy: Frequency Autoregressive Visuomotor Policy with Continuous Tokens

### v1

@@ 1 Introduction
Figure 1: Action signals reconstructed in different frequency domain ranges.
@@ 2.2 Visuomotor Policy for Robotic Manipulation
Figure 2: Pipeline of FreqPolicy showing both training (a) and inference (b) procedures. We first transforms action trajectories into the frequency domain via DCT, and then learns latent codes for different frequency level actions using FreqPolicy, and reconstructs actions through masked prediction and a diffusion-based decoder. This enables robust, frequency-aware, and high-fidelity robotic action generation.
@@ 3.1 Overview
Problem Formulation Given a dataset of paired sequences $\mathcal{D}=(\mathbf{o},\mathbf{x})$, where each $\mathbf{o}=[o_{1},\ldots,o_{N}]^{T}$ denotes a sequence of observations (e.g., RGB images, depth maps, or point clouds) and each $\mathbf{x}=[x_{1},x_{2},\ldots,x_{N}]^{T}$ represents the corresponding sequence of robot actions, our objective is to train a policy that can produce action trajectories $\hat{\mathbf{x}}$ based on new observation sequences $\hat{\mathbf{o}}$ to effectively perform the demonstrated tasks.
As shown in Figure 2, our method first transforms action trajectories into the frequency domain using the Discrete Cosine Transform (DCT), decomposing them into frequency components. These trajectories are then recovered by Inverse DCT but mask partial frequency components, and a FreqPolicy encoder maps the recovered trajectories, along with conditional embeddings and index information, into latent codes corresponding to different frequency levels. Then the FreqPolicy decoder reconstructs the masked frequency components from their latent representations, ensuring that key frequency components are selectively preserved or predicted. To further enhance generative modeling, we introduce a diffusion-based decoder that takes as input the trajectory latent code, frequency index, and condition embeddings, and generates refined trajectories by minimizing diffusion loss. This approach enables the model to learn robust, frequency-aware representations and generate high-fidelity robotic actions by leveraging both global structure and fine-grained motion details across different frequency domains. This section is organized as follows. We begin with an analysis of action trajectories in the frequency domain, followed by an introduction to the space-frequency transformation method employed in our approach. Then, we describe how to integrate frequency-aware action encoding with continuous tokens. Finally, we present FreqPolicy, a novel frequency-domain autoregressive framework illustrated in Figure 2.
@@ 3.2 Frequency Domain Analysis
Performance under Frequency Compression. To further assess the role of high-frequency information on task performance across different benchmarks, we compress action signals by removing the high-frequency components from the original sequences, reconstructed the actions with varying proportions of low-frequency information, and measured their success rates on the original tasks. The results, shown in Figure 3(b), indicate that omitting high-frequency signals leads to a significant drop in success rates for Adroit tasks. In contrast, low-dimensional tasks in Robomimic maintain stable performance with as little as 12% of low-frequency information preserved. This analysis underscore the importance of frequency-aware modeling: the necessity of high-frequency signals for effective action representation varies across different scenarios.
@@ 3.3 DCT Decomposition
Our spectral analysis reveals that the proportion of frequency components required for effective action representation varies across different tasks. Therefore, it is necessary to transform the action space from the time domain to the frequency domain, for more effective robotic action modeling. To efficiently achieve this, we employ the Discrete Cosine Transform (DCT), which can project time-domain trajectories onto a set of cosine basis functions with low computational cost. This provides a compact and interpretable frequency-domain representation, enabling hierarchical modeling of actions in various tasks settings and scenarios. Specifically, given an action sequence $\mathbf{x}=[x_{1},x_{2},\ldots,x_{N}]^{T}$ into the frequency domain, its DCT is defined as:
@@
$y_{n}^{k}=\frac{1}{N}\left[\frac{1}{2}X_{0}+\sum_{i=1}^{k-1}X_{i}\cos\left(\frac{\pi}{N}\left(n+\frac{1}{2}\right)i\right)\right],\quad n=0,1,\ldots,N-1$ (2)
Specifically, let $\mathbf{y}^{k}=[y_{0}^{k},y_{2}^{k},\ldots,y_{N-1}^{k}]^{T}$ denote the $k$-level reconstruction of the original sequence $\mathbf{x}$. If $k=N$, meaning all DCT coefficients are used, the reconstruction will recover the original trajectory without any loss of information. However, if $k<N$, only the lowest $k$ frequency components are kept, and all higher-frequency components are discarded. As a result, the reconstructed trajectory $\mathbf{y}^{k}$ becomes a compressed and smoother version of the original, preserving the essential structure of the original sequence while discarding minor fluctuations.
@@ 3.5 Frequency-based Hierarchical Generation
Our previous analysis demonstrated that compressing actions in the frequency domain at different scales reveals hierarchical information. Actions reconstructed with higher compression ratio are smoother and better capture the overall trends of the motion, which helps reduce noise interference and lowers the difficulty of generation. Based on this observation, we propose a frequency-based autoregressive strategy for hierarchical action generation. As illustrated in Figure 2. Starting from zero input, we iterate from low to high frequencies $N_{\text{iter}}$ times to generate the action sequence. In each iter $i\in\{1,2,...,N_{\text{iter}}\}$, the observation tokens, the frequency level index $l_{i-1}\in\{l_{0},l_{1},...,l_{N_{\text{iter}}}\}$, and the corresponding $l_{i-1}$ level reconstruction action tokens $\mathbf{y}^{l_{i-1}}$ are encoded to produce a continuous token $\mathbf{z}^{l_{i-1}}$ that represents the current context. The diffusion model then samples a full-frequency action sequence $\hat{\mathbf{x}}^{i}$ conditioned on this token. For the next step, we use the $l_{i}$-level reconstruction of $\hat{\mathbf{x}}^{i}$, i.e., $\mathbf{y}^{l_{i}}$, as the input for the following step, which contains richer information than previous input.
@@ 4 Experiments
Table 1: Comparison on 10 simulation tasks in Adroit, DexArt and Meta-World. To ensure fair comparison, all tasks employ full-spectrum outputs and identical frequency level progression. * denotes results we reproduced using the same expert demonstrations as ours.
Adroit [30] DexArt [4] Meta-World [46] Average
[…]
Mamba Policy* $\mathbf{100\scriptstyle{\pm 0}}$ $59\scriptstyle{\pm 3}$ $55\scriptstyle{\pm 2}$ $79\scriptstyle{\pm 3}$ $\mathbf{35\scriptstyle{\pm 6}}$ $65\scriptstyle{\pm 5}$ $23\scriptstyle{\pm 2}$ $96\scriptstyle{\pm 2}$ $90\scriptstyle{\pm 2}$ $82\scriptstyle{\pm 5}$ 68.4
@@ 4.1 Experimental Setup
4.2 Simulation Result
High Success Rate. In Table 1, we compare our method on 10 tasks from Adroit, DexArt, and MetaWorld, while Table 3 presents results on 7 RoboTwin tasks using colored point clouds as observations. In Table 1, our method outperforms both the state-of-the-art (SOTA) methods and a baseline (our codebase without DCT composition) on 8 tasks, and achieves comparable results on the remaining 2. Similarly, as shown in Table 3, our method achieves superior performance in RoboTwin, with 5 tasks reaching higher success rates than DP and DP3. We also evaluated all 48 tasks in Adroit, DexArt, and MetaWorld and reported the average success rate for each benchmark. As shown in Table 2), our approach maintains an average success rate of 67.9%, surpassing Diffusion Policy 3D (64.6%) and Mamba Policy (65.8%). These results demonstrate the effectiveness of our approach across a variety of task settings and scenarios. More detailed results for each individual task are provided in the appendix.
Generalization Ability Results in Table 4 demonstrate that our method achieves consistently strong performance in unseen DexArt test environments, maintaining high success rates under both limited (10 demonstrations) and abundant (100 demonstrations) data settings. These results demonstrate the method’s robustness and generalization ability in diverse scenarios.
Discrete vs Continuous As shown in Table 5, our systematic comparison between discrete autoregressive and continuous diffusion models reveals a key finding: discrete representation methods (CARP, BeT) exhibit significant limitations in modeling continuous action spaces, primarily due to information loss during the discretization process, particularly evident in hard tasks requiring precise control. In contrast, our proposed approach that integrates autoregressive and diffusion mechanisms in a continuous representation framework not only maintains consistent high performance across various tasks, but also inherits the computational efficiency advantages of autoregressive methods (requiring only 1/10 of DP’s inference time), conclusively demonstrating its superior expressiveness, computational efficiency, and environmental adaptability in modeling continuous action spaces.
Table 2: Main results on 48 simulation tasks. Averaged over tasks,* denotes results we reproduced using the same expert demonstrations, for fair comparison.Success rates for individual tasks are in Appendix.
@@ 4.2 Simulation Result
[…]
@@
Table 3: Comparison on the RoboTwin Benchmark for Dual-Arm Manipulation with D435 Camera Setting. We evaluated our approach on 7 tasks using 20 expert demonstrations and 3 seeds (0, 1, 2) and report the success rate. Our method was compared against DP3 (XYZ+RGB) and DP, all tested under the same conditions. Both DP and our method are trained for 500 epochs, while DP3 is trained for 3,000 epochs.
@@
DP $0.0\scriptstyle{\pm 0.0}$ $0.0\scriptstyle{\pm 0.0}$ $6.3\scriptstyle{\pm 5.9}$ $1.7\scriptstyle{\pm 0.6}$ $0.0\scriptstyle{\pm 0.0}$ $5.3\scriptstyle{\pm 2.5}$ $8.0\scriptstyle{\pm 2.0}$
@@
Table 4: Generalization Results. Comparison on unseen DexArt test data after 3000 training epochs, averaged over 100 trials $\times$ 3 random seeds.
[…]
@@
DP3 $12.2\scriptstyle{\pm 2}$ $11.5\scriptstyle{\pm 3}$ $8.5\scriptstyle{\pm 1}$ $15.0\scriptstyle{\pm 3}$ $36.4\scriptstyle{\pm 5}$ $\mathbf{17.4\scriptstyle{\pm 2}}$ $25.4\scriptstyle{\pm 1}$ $\mathbf{29.2\scriptstyle{\pm 6}}$
@@
Table 5: Comparisons on Robomimic benchmark [24] and Push-T task [12].

### v2

@@ 1 Introduction
Figure 1: Action signals reconstructed across different frequency bands; see Appendix for energy details.
@@ 2.2 Visuomotor Policy for Robotic Manipulation
Figure 2: Pipeline of FreqPolicy showing both training (a) and inference (b) procedures. We first transform action trajectories into the frequency domain via DCT, and then learns latent codes for different frequency level actions using FreqPolicy, and reconstructs actions through masked prediction and a diffusion-based decoder. This enables robust, frequency-aware, and high-fidelity robotic action generation.
@@ 3.1 Overview
Problem Formulation Given a dataset of paired sequences $\mathcal{D}=\{(\mathbf{o},\mathbf{x})\}$, where each $\mathbf{o}=[o_{0},\ldots,o_{N-1}]^{T}$ denotes a sequence of observations (e.g., RGB images, depth maps, or point clouds) and each $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N-1}]^{T}\in\mathbb{R}^{N\times d}$ represents the corresponding sequence of robot actions with variable length $N$ and action dimension $d$, our objective is to train a policy that can produce action trajectories $\hat{\mathbf{x}}$ based on new observation sequences $\hat{\mathbf{o}}$ to effectively perform the demonstrated tasks.
As shown in Figure 2, our method consists of two stages. During training (Figure 2a), we apply DCT to decompose action sequences into frequency components and reconstruct them at different levels via inverse DCT. For frequency level $k$, the FreqPolicy encoder-decoder processes the $k$-level reconstruction $\mathbf{y}^{k}$, observation features, and frequency index to produce a continuous token $\mathbf{z}^{k}$, which conditions a diffusion model to reconstruct the original full-frequency action $\mathbf{x}$. We adopt frequency-aware masking (gray blocks in the figure) to improve training efficiency. During inference (Figure 2b), we perform hierarchical generation through $N_{\text{iter}}$ iterations. Starting from zero ($\mathbf{y}^{l_{0}}=\mathbf{0}$), each iteration $i$ generates a full-frequency action $\hat{\mathbf{x}}^{i}$ conditioned on the previous output $\mathbf{y}^{l_{i-1}}$, then filters it to $\mathbf{y}^{l_{i}}$ for the next iteration. This coarse-to-fine process progressively refines actions from low-frequency global structure to high-frequency fine details.
This section is organized as follows: Section 3.2 analyzes action trajectories in the frequency domain; Section 3.3 introduces the space-frequency transformation method employed in our approach; Section 3.4 describes the continuous token representation with diffusion models; and Section 3.5 presents the frequency-based hierarchical generation strategy.
@@ 3.2 Frequency Domain Analysis
Performance under Frequency Compression. To further assess the role of high-frequency information on task performance across different benchmarks, we compress action signals by removing the high-frequency components from the original sequences, reconstructed the actions with varying proportions of low-frequency information, and measured their success rates on the original tasks. The results, shown in Figure 3(b), indicate that omitting high-frequency signals leads to a significant drop in success rates for Adroit tasks. In contrast, low-dimensional tasks in Robomimic maintain stable performance with as little as 12% of low-frequency information preserved. This analysis underscores the importance of frequency-aware modeling: the necessity of high-frequency signals for effective action representation varies across different scenarios.
@@ 3.3 DCT Decomposition
Our spectral analysis reveals that the proportion of frequency components required for effective action representation varies across different tasks. Therefore, it is necessary to transform the action space from the time domain to the frequency domain, for more effective robotic action modeling. To efficiently achieve this, we employ the Discrete Cosine Transform (DCT), which can project time-domain trajectories onto a set of cosine basis functions with low computational cost. This provides a compact and interpretable frequency-domain representation, enabling hierarchical modeling of actions in various tasks settings and scenarios. Specifically, given an action sequence $\mathbf{x}=[x_{0},x_{1},\ldots,x_{N-1}]^{T}\in\mathbb{R}^{N\times d}$, where each $x_{n}\in\mathbb{R}^{d}$ represents the action at time step $n$, we transform it into the frequency domain by applying the DCT independently to each action dimension. For each dimension, the DCT is defined as:
@@
$y_{n}^{k}=\frac{1}{N}\left[X_{0}+2\sum_{i=1}^{k-1}X_{i}\cos\left(\frac{\pi}{N}\left(n+\frac{1}{2}\right)i\right)\right],\quad n=0,1,\ldots,N-1$ (2)
Specifically, let $\mathbf{y}^{k}=[y_{0}^{k},y_{1}^{k},\ldots,y_{N-1}^{k}]^{T}\in\mathbb{R}^{N\times d}$ denote the $k$-level reconstruction of the original sequence $\mathbf{x}$. If $k=N$, meaning all DCT coefficients are used, the reconstruction will recover the original trajectory without any loss of information. However, if $k<N$, only the lowest $k$ frequency components are kept, and all higher-frequency components are discarded. As a result, the reconstructed trajectory $\mathbf{y}^{k}$ becomes a compressed and smoother version of the original, preserving the essential structure of the original sequence while discarding minor fluctuations.
@@ 3.5 Frequency-based Hierarchical Generation
Our previous analysis demonstrated that compressing actions in the frequency domain at different scales reveals hierarchical information. Actions reconstructed with higher compression ratio are smoother and better capture the overall trends of the motion, which helps reduce noise interference and lowers the difficulty of generation. Based on this observation, we propose a frequency-based autoregressive strategy for hierarchical action generation, as illustrated in Figure 2. We define an increasing frequency level sequence $\{l_{0},l_{1},\ldots,l_{N_{\text{iter}}}\}$ where $l_{0}=0$ (starting from zero input) and $l_{N_{\text{iter}}}=N$ (full frequency). Each $l_{i}$ represents the number of DCT coefficients retained at iteration $i$. Starting from $l_{0}$, we iterate from low to high frequencies $N_{\text{iter}}$ times to generate the action sequence. In each iteration $i\in\{1,2,...,N_{\text{iter}}\}$, the observation tokens, the frequency level index $l_{i-1}$, and the corresponding $l_{i-1}$-level reconstruction action tokens $\mathbf{y}^{l_{i-1}}$ are encoded to produce a continuous token $\mathbf{z}^{l_{i-1}}$ that represents the current context. The diffusion model then samples a full-frequency action sequence $\hat{\mathbf{x}}^{i}$ conditioned on this token. For the next step, we use the $l_{i}$-level reconstruction of $\hat{\mathbf{x}}^{i}$, i.e., $\mathbf{y}^{l_{i}}$, as the input for the following iteration, which contains richer frequency information than the previous input.
Example. For a 16-timestep trajectory with $N_{\text{iter}}=4$ iterations, we might use $\{l_{0}=0,l_{1}=4,l_{2}=8,l_{3}=12,l_{4}=16\}$. At iteration 1, we start with $\mathbf{y}^{0}$ (all zeros) and generate $\hat{\mathbf{x}}^{1}$, then filter it to keep only the first 4 DCT coefficients, obtaining $\mathbf{y}^{4}$ for iteration 2. This process continues, progressively adding more frequency components ($4\to 8\to 12\to 16$) until the final full-frequency output.
@@ 4.1 Experimental Setup
Table 1: Success rate (%) comparison on 10 simulation tasks in Adroit, DexArt and Meta-World. To ensure fair comparison, all tasks employ full-spectrum outputs and identical frequency level progression. * denotes results we reproduced using the same expert demonstrations as ours.
Adroit [30] DexArt [4] Meta-World [46] Average
Alg $\backslash$ Task Hammer Door Pen Laptop Faucet Toilet Bucket Assembly Disassemble Stick Push
Mamba Policy* $\mathbf{100\scriptstyle{\pm 0}}$ $9\scriptstyle{\pm 3} 55$$5\scriptstyle{\pm 2}$ $79\scriptstyle{\pm 3}$ $\mathbf{35\scriptstyle{\pm 6}}$ $65\scriptstyle{\pm 5}$ $23\scriptstyle{\pm 2}$ $96\scriptstyle{\pm 2}$ $90\scriptstyle{\pm 2}$ $82\scriptstyle{\pm 5}$ 68.4
Table 2: Main results on 48 simulation tasks. The table reports average success rates (%) across all tasks. * denotes results we reproduced using the same expert demonstrations for fair comparison. Success rates for individual tasks are in Appendix.
@@
Algorithm $\backslash$ Task (3) (4) Easy (20) Medium (11) Hard (5) Very Hard (5)
@@
Table 3: Success rate (%) comparison on the RoboTwin Benchmark for Dual-Arm Manipulation with D435 Camera Setting. We evaluated our approach on 7 tasks using 20 expert demonstrations and 3 seeds (0, 1, 2) and report the success rate. Our method was compared against DP3 (XYZ+RGB) and DP, all tested under the same conditions. Both DP and our method are trained for 500 epochs, while DP3 is trained for 3,000 epochs.
@@
DP $\scriptstyle{\pm 0.0} 0.0±0.0$$\scriptstyle{\pm 0.0}$ $6.3\scriptstyle{\pm 5.9}$ $1.7\scriptstyle{\pm 0.6}$ $0.0\scriptstyle{\pm 0.0}$ $5.3\scriptstyle{\pm 2.5}$ $8.0\scriptstyle{\pm 2.0}$
@@
Table 4: Generalization Results. Success rates (%) on unseen DexArt test data after 3000 training epochs, averaged over 100 trials $\times$ 3 random seeds.
Alg $\backslash$ Task 10 Demonstrations 100 Demonstrations
@@
DP3 $2.2\scriptstyle{\pm 2} 11$$1.5\scriptstyle{\pm 3}$ $8.5\scriptstyle{\pm 1}$ $15.0\scriptstyle{\pm 3}$ $36.4\scriptstyle{\pm 5}$ $\mathbf{17.4\scriptstyle{\pm 2}}$ $25.4\scriptstyle{\pm 1}$ $\mathbf{29.2\scriptstyle{\pm 6}}$
@@
Table 5: Noise Robustness Evaluation. Success rates (%) with relative change vs. high-quality baseline (see Appendix Table 10) shown as superscripts. We evaluate two noise conditions: (1) Low-Quality demonstrations using suboptimal trajectories (5 tasks, hammer excluded), and (2) Gaussian Noise (std=0.025, 0.05, 0.1) applied to high-quality baseline demonstrations (6 tasks).
Method Low-Quality Demonstrations Gaussian Noise std=0.025
Pick Soccer Stick Hammer Door Pen Avg. Pick Soccer Stick Hammer Door Pen Avg.
DP3 $10^{-17}$ $15^{-32}$ $41^{-28}$ $47^{-11}$ $33^{-34}$ $29^{-40}$ $26^{+117}$ $25^{+14}$ $\mathbf{63^{+11}}$ $0^{-100}$ $\mathbf{55^{+4}}$ $53^{+6}$ $37^{-24}$
Mamba $16^{-36}$ $\mathbf{20^{-29}}$ $0^{-27} 44$$4^{-25}$ $\mathbf{34^{-38}}$ $31^{-43}$ $18^{-28}$ $\mathbf{33^{+18}}$ $61^{+11}$ $0^{-100}$ $50^{-15}$ $\mathbf{57^{+4}}$ $37^{-32}$
Ours $\mathbf{20^{-33}}$ $17^{-47}$ $\mathbf{43^{-31}}$ $\mathbf{49^{-25}}$ $32^{-45}$ $\mathbf{32^{-44}}$ $\mathbf{32^{+7}}$ $\mathbf{33^{+3}}$ $60^{-3}$ $\mathbf{70^{-30}}$ $52^{-20}$ $56^{-3}$ $\mathbf{51^{-13}}$
Method Gaussian Noise std=0.05 Gaussian Noise std=0.1
Pick Soccer Stick Hammer Door Pen Avg. Pick Soccer Stick Hammer Door Pen Avg.
DP3 $21^{+75}$ $19^{-14}$ $51^{-11}$ $0^{-100}$ $27^{-49}$ $55^{+10}$ $29^{-41}$ $7^{-42}$ $18^{-18}$ $50^{-12}$ $0^{-100}$ $24^{-55}$ $38^{-24}$ $23^{-53}$
Mamba $23^{-8}$ $25^{-11}$ $\mathbf{60^{+9}}$ $0^{-100}$ $\mathbf{51^{-14}}$ $\mathbf{57^{+4}}$ $36^{-33}$ $3^{-88}$ $14^{-50}$ $39^{-29}$ $0^{-100}$ $21^{-64}$ $39^{-29}$ $19^{-64}$
Ours $\mathbf{25^{-17}}$ $\mathbf{30^{-6}}$ $\mathbf{60^{-3}}$ $\mathbf{51^{-49}}$ $46^{-29}$ $55^{-5}$ $\mathbf{45^{-23}}$ $\mathbf{20^{-33}}$ $\mathbf{27^{-16}}$ $\mathbf{52^{-16}}$ $\mathbf{15^{-85}}$ $\mathbf{32^{-51}}$ $\mathbf{49^{-16}}$ $\mathbf{33^{-44}}$
4.2 Simulation Result
High Success Rate. In Table 1, we compare our method on 10 tasks from Adroit, DexArt, and MetaWorld, while Table 3 presents results on 7 RoboTwin tasks using colored point clouds as observations. In Table 1, our method outperforms both the state-of-the-art (SOTA) methods and a baseline (without DCT decomposition) on 8 tasks, and achieves comparable results on the remaining 2. Similarly, as shown in Table 3, our method achieves superior performance in RoboTwin, with 5 tasks reaching higher success rates than DP and DP3. We also evaluated all 48 tasks in Adroit, DexArt, and MetaWorld and reported the average success rate for each benchmark. As shown in Table 2, our approach maintains an average success rate of 67.9%, surpassing Diffusion Policy 3D (64.6%) and Mamba Policy (65.8%). These results demonstrate the effectiveness of our approach across a variety of task settings and scenarios. More detailed results for each individual task are provided in the appendix.
Generalization Ability. Results in Table 4 demonstrate that our method achieves consistently strong performance in unseen DexArt test environments, maintaining high success rates under both limited (10 demonstrations) and abundant (100 demonstrations) data settings. These results demonstrate the method’s robustness and generalization ability in diverse scenarios.
Noise Robustness. We evaluate robustness under two noise scenarios (Table 5): (1) training on low-quality demonstrations with reduced reward thresholds (5 tasks used, hammer task cannot generate low-reward expert demonstrations), and (2) adding Gaussian noise (std=0.025, 0.05, 0.1) to high-quality demonstrations. On low-quality demonstrations, all methods show significant performance decline, validating the importance of high-quality training data. More importantly, under Gaussian noise conditions, our method demonstrates exceptional robustness: at high noise (std=0.1), we achieve 33% success versus 19-23% for baselines. The Hammer task is particularly revealing—baselines completely fail (0% success) under any Gaussian noise, while we maintain 70%, 51%, and 15% at increasing noise levels. This superior performance stems from our low-frequency to high-frequency progressive generation mechanism’s natural anti-noise advantages: (1) Low-frequency anti-noise characteristics—low-frequency components inherently have stronger noise resistance, establishing a stable action foundation in initial stages; (2) Progressive filtering—each frequency domain level’s reconstruction process acts as natural noise filtering; (3) Structured constraints—the coarse-to-fine generation process ensures that key action structures are preserved even under noise interference. Interestingly, moderate noise (std=0.025) can promote generalization in certain tasks: for Pen spinning and Pick Out of Hole, low-intensity Gaussian noise unintentionally increases hand movement variation, adding action diversity and leading to higher success rates than noise-free baselines. This suggests beneficial noise augmentation can enrich the action distribution.
Table 6: Success rates (%) comparisons on Robomimic benchmark [24] and Push-T task [12].